OpenStudy (anonymous):

which is a solution of the equation y = 2x + 5 (5, 20) (6, 12) (3, 11) (4, 14)

OpenStudy (anonymous):

@texaschic101

OpenStudy (arabpride):

OK simple, when u look at the coordinate points, do u know which one iz the 'x' value nd which one ix the 'y' value??

OpenStudy (arabpride):

@allopersonwhat

OpenStudy (anonymous):

yah

OpenStudy (texaschic101):

sub your answer choices into the problem and see which one is true

OpenStudy (anonymous):

but like how do I solve it that's where im confused

OpenStudy (arabpride):

all you have to do iz plug the valuez into the equation, and see which one comes out equal..

OpenStudy (anonymous):

ok thx :)

OpenStudy (anonymous):

now I don't know who to give a medal ahhh

OpenStudy (anonymous):

ok this is what im gonna do im going to fan you arab and medal texas cuz im already fanning her.

OpenStudy (arabpride):

y = 2x + 5 (5, 20) (6, 12) (3, 11) (4, 14) 5,20 : (20) = 2(5) + 5 20 = 10 + 5 20 does not equal 15 so its not (5,20) 6,12 : (12) = 2(6) + 5 12= 12 + 5 12 does not equal 17 so its not (6,12) 3,11 : (11) = 2(3) + 5 11 = 6 + 5 11 = 11 SO yes it izz (3,11) 4,14 : (14) = 2(4) + 5 14 = 8 + 5 14 does not equal 13 so it is not (4,14) @allopersonwhat

OpenStudy (anonymous):

oh woooooo thank you so much that's awesome :)

OpenStudy (arabpride):

You're Welcome No Problem, sorry i took a while..